OpenStudy (anonymous):

Is square root 1 minus sine squared theta = cos Θ true? If so, in which quadrants does angle Θ terminate?

OpenStudy (anonymous):

lyn can u help me next :3

OpenStudy (anonymous):

False True; quadrants I & IV True; quadrants II & III True; quadrants I & III

OpenStudy (lynfran):

\[\cos ^{2}\theta+\sin ^{2}\theta=1\]\[\cos ^{2}\theta=1-\sin ^{2}\theta \]\[\cos \theta=\sqrt{(1-\sin ^{2} \theta)}\]

OpenStudy (anonymous):

how do i know which quadrant its in?

OpenStudy (anonymous):

@LynFran ^^^

OpenStudy (lynfran):

the 1st and last quadrant

OpenStudy (lynfran):

thats where cos is positive ...

OpenStudy (anonymous):

ohhh ok i see. thanks so so much. hey do you ming helping me with a couple more? you don't have to if you don't want. @LynFran

OpenStudy (lynfran):

so tag me in new post